# Coordinating Search-Informed Reasoning and Reasoning-Guided Search in Claim Verification

### v1

@@ Abstract
Multi-hop claim verification is inherently challenging, requiring multi-step reasoning to construct verification chains while iteratively searching for information to uncover hidden bridging facts. This process is fundamentally interleaved, as effective reasoning relies on dynamically retrieved evidence, while effective search demands reasoning to refine queries based on partial information. To achieve this, we propose Hierarchical Agent Reasoning and Information Search (HARIS), explicitly modeling the coordinated process of reasoning-driven searching and search-informed reasoning. HARIS consists of a high-level reasoning agent that focuses on constructing the main verification chain, generating factual questions when more information is needed, and a low-level search agent that iteratively retrieves more information, refining its search based on intermediate findings. This design allows each agent to specialize in its respective task, enhancing verification accuracy and interpretability. HARIS is trained using reinforcement learning with outcome-based rewards. Experimental results on the EX-FEVER and HOVER benchmarks demonstrate that HARIS achieves strong performance, greatly advancing multi-hop claim verification.
@@ 1 Introduction
Claim verification (Guo et al., 2022) has become a critical challenge as misinformation proliferates online. It requires systems to determine whether a given claim is supported or refuted based on retrieved evidence. While verifying simple claims involves shallow reasoning within a single document, the verification of complex, multi-hop claims presents a fundamentally different challenge. This difficulty stems from the fragmented nature of evidence (Pham et al., 2025; Atanasova et al., 2022). Effectively verifying such claims requires an joint process of multi-step reasoning and iterative information searching (Zheng et al., 2025).
Figure 1: Example of multi-hop verification. Verifying the claim requires coordinating reasoning-guided search to disambiguate entities and search-informed reasoning to adapt based on retrieved evidence. The correct path—from Baby (2007) to Overwatch via Feodor Chin—emerges only through this dynamic interplay, not static decomposition or single-pass retrieval.
A key challenge in multi-hop verification is identifying the correct bridging facts—implicit links that connect separate pieces of evidence but are not explicitly stated. As illustrated in Figure 1, verifying this claim requires identifying Feodor Chin as the critical bridging fact between the film Baby (2007) and the character Zenyatta from the game Overwatch, among other irrelevant paths. This process necessitates intensive interactions between reasoning, which proposes candidate hypotheses to identify potential bridging facts, and iterative search, which retrieves evidence to validate or eliminate certain hypotheses. In particular, reasoning is necessary to construct verification chains, but effective reasoning depends on relevant evidence, which often requires iterative search. Meanwhile, effective search relies on reasoning to formulate queries based on partially retrieved evidence. This reciprocal relationship, where reasoning shapes search and retrieved evidence refines ongoing reasoning, captures the recursive nature of multi-hop verification.
@@
To tackle these challenges, we propose HARIS—Hierarchical Agent Reasoning and Information Search, explicitly modeling the coordinated process of reasoning-driven searching and search-informed reasoning. HARIS consists of two specialized large language model (LLM) agents: a high-level reasoning agent and a low-level search agent, both trained using reinforcement learning to optimize their respective tasks. The high-level agent forms the main verification chain, generating factual questions when additional information is needed. The low-level agent handles these questions through dynamic search, iteratively refining its queries based on partial results to progressively build a comprehensive evidence base. This design allows each agent to specialize in its respective task, enhancing both verification accuracy and interpretability by clearly modeling the mutually reinforcing interaction between reasoning and information searching.
@@
  • HARIS is trained end-to-end via Group Relative Policy Optimization (GRPO) using outcome-based rewards, directly optimizing task performance without intermediate supervision.
@@ 2.1 Claim Verification
Claim verification research has been increasingly focusing on enhancing transparency during decision making (Zeng and Gao, 2024; Chen et al., 2024; Hu et al., 2025a). QACheck (Pan et al., 2023a) reformulates verification as a progressive question-answering (QA) task, validating claims through step-wise questioning. Structural approaches (Jeon and Lee, 2025; Pham et al., 2025), such as ProgramFC (Pan et al., 2023b) and FOLK (Wang and Shu, 2023), which use symbolic reasoning or reasoning program to enforce systematic verification. The Decompose-Then-Verify (Wanner et al., 2024b, a; Jiang et al., 2024; Hu et al., 2025b; Lu et al., 2025) paradigm focus on breaking down complex claims into simpler sub-claims for independent validation. Agentic approaches have also emerged as a promising direction (Zhao et al., 2024). For example, LoCal (Ma et al., 2025) introduces an in-context learning-based multi-agent framework that emphasizes logical and causal consistency. These approaches collectively aim to provide clearer insights into the reasoning logic behind model predictions.
@@ 2.2 Reasoning & Searching
Recent work has greatly expanded the reasoning capabilities of LLMs by integrating external search mechanisms (Xiong et al., 2025; Guan et al., 2025; Sun et al., 2025). Notable methods like Search-o1 (Li et al., 2025) incorporate dynamic search into reasoning frameworks, improving factual accuracy in open-domain and multi-hop reasoning tasks leveraging large reasoning models. Furthermore, reinforcement learning (RL) methods like Group Relative Policy Optimization (GRPO) (Shao et al., 2024) have been used to incentivize search capabilities in LLMs (Gao et al., 2025), encouraging models to autonomously generate context-aware queries and effectively integrate retrieved information (Jin et al., 2025; Song et al., 2025; Chen et al., 2025; Qian et al., 2025; Wu et al., 2025). Collectively, these approaches demonstrate that RL-trained search integration can effectively improve performance on knowledge-intensive tasks.
@@ 3.1 Why Reasoning and Search Agents?
Meanwhile, search and reasoning agents interact closely to mutually enhance their performance. This mutual enhancement is realized through both interactive generation, where agents iteratively guide each other’s outputs, and coordinated training, where the agents are alternately optimized to reinforce their collaborative capabilities, as illustrated in Figure 2.
@@ 3.2 High-level Reasoning Agent
The high-level reasoning agent in HARIS constructs the main claim verification chain, coordinating multi-step reasoning and generating factual questions for the search agent when additional information is needed.
@@ 3.2.1 Reasoning Agent Rollout
Following prior work (Chen et al., 2025; Jin et al., 2025), the reasoning agent’s rollout process is guided by special tags that define question actions. Specifically, the tags <question> and </question> indicate that an action to call the search agent should be invoked. When </question> tag is identified, the generation is paused and the content wrapped within will be regarded as the factual question and sent to the search agent. The reported information from search agent is wrapped within <result> tags and appended to the sequence to enable continued rollout. The rollout ends when a final verification is derived, wrapped within <verification> tags. The prompt template is provided in Appendix A.6.
@@ 3.2.2 Reasoning Agent Reward
The reasoning agent is responsible for performing multi-step reasoning over the gathered evidence to verify the original claim. Its reward is the correctness of the final verification decision. Formally, the overall reward $R_{\text{high}}$ is given by:
@@ 3.3.1 Training Data Synthesis
To train the search agent, it is crucial that the training data not only includes diverse question-answer pairs but also closely aligns with the style and distribution of the questions generated by the high-level agent. This alignment ensures that the search agent can effectively collaborate with the reasoning agent during multi-hop verification. To create such training set, we sample questions from the high-level agent rollouts, pair them with the corresponding ground-truth evidence, and use GPT-4o (OpenAI, 2024b) to generate pseudo ground-truth answers. To maintain data quality, we filter out pairs where GPT-4o outputs "none" as the answer, typically removing about 10% of the data. This process synthesize data that is both contextually relevant and closely matched to the reasoning patterns of the high-level agent. Synthesis details and examples can be found in Appendix A.4 and Table B.1.
@@ 3.3.2 Search Agent Rollout
The search agent rollout process is similar to the reasoning agent. When the generation process encounters a </search> tag, it pauses and extracts the content wrapped within the tags as the search query. This query is then used to perform top-k retrieval from the knowledge corpus. The retrieved text is wrapped with <result> and </result> tags and appended to the paused sequence, allowing the rollout to continue iteratively until the agent determines that sufficient information has been gathered. At this point, the agent reports the collected evidence within <report> and </report> tags, marking the completion of the rollout. The prompt template can be found in Appendix A.6.
@@ 3.3.3 Search Agent Reward
We use a combination of format and LLM-as-a-Judge approaches for search agent rewards. The LLM-as-a-Judge approach evaluates the quality the gathered information using an LLM, comparing the final output against the pseudo ground-truth answer [^1] We use GPT-4o-mini(OpenAI, 2024a) for this purpose, as it is cost-efficient. The evaluation score is computed as:
@@
where ${a_{\text{pred}}}$ is the search agent’s final output and ${a_{\text{gt}}}$ is the pseudo ground-truth answer. This approach is less strict than exact match (EM) metrics, better aligning with the search agent’s goal of collecting comprehensive, contextually relevant information rather than just short, exact answers. The implementation can be found in Appendix A.5.1. The formatting reward is set to 0.1, following Chen et al. (2025). Formally, the reward $R_{\text{low}}$ is given by:
@@ 3.5 Coordinated Agent Training
To enable effective collaboration between the high-level reasoning agent and the low-level search agent, we adopt a Coordinated Agent Training strategy. This training scheme allows for both agents to evolve in tandem, maintaining alignment while progressively refining their specialized capabilities through mutual interaction.
Specifically, we use a training set $T$ containing claim verification data and train the agents over two stages. In the initial stage, the untrained low-level search agent $L_{0}$ and high-level reasoning agent $H_{0}$ are trained sequentially to establish their foundational capabilities. First, we sample questions from $H_{0}$ to train $L_{0}$, producing the updated search agent $L_{1}$. This updated search agent then supports the training of $H_{0}$, producing the updated reasoning agent $H_{1}$. This stage establishes a foundation for both agents in reasoning, searching, and formatting skills.
In the second stage, we alternate training between the two agents to promote coordination. The dataset $T$ is divided into $N$ segments. For each segment, we sample questions from the current high-level agent $H_{i}$ to train the low-level agent, producing $L_{i+1}$. The updated low-level agent $L_{i+1}$ is then used to colloraborate the high-level agent, resulting in $H_{i+1}$. This alternating process continues for $N$ rounds, fostering mutual adaptation and ensuring the agents remain closely aligned throughout training.
@@
Input: Initial low-level agent $L_{0}$ and high-level agent $H_{0}$, training set $T$
@@
for epoch $=1$ to $2$ do
     if epoch $=1$ then
@@
     else
         Divide $T$ into $N$ segments
         for each segment $T_{i}\in T$ do
@@
RAG $58.96$ $59.20$ $56.59$ $56.60$ $55.06$ $55.20$ $68.83$ $69.00$ $64.41$ $64.80$ $56.40$
Decomp-Verify $62.39$ $62.60$ $57.27$ $57.31$ $54.55$ $55.60$ $68.03$ $68.40$ $61.95$ $63.00$ $36.00$
ProgramFC $66.84$ $66.80$ $55.35$ $56.80$ $48.30$ $52.60$ $71.28$ $71.60$ $60.34$ $62.40$ $21.60$
QACheck $67.60$ $67.60$ $60.60$ $60.60$ $58.91$ $59.00$ $75.52$ $75.60$ $68.42$ $68.60$ $58.00$
FOLK $67.22$ $67.60$ $59.89$ $61.20$ $50.90$ $55.20$ $75.55$ $75.80$ $67.24$ $68.40$ $50.70$
Search-o1 $68.72$ $69.00$ $59.34$ $59.80$ $54.90$ $56.60$ $77.41$ $77.80$ $72.08$ $72.80$ $52.40$
@@ 4.1 Datasets
For training, we sample 7,200 data from the combined training data of EX-FEVER and HOVER. For evaluation, following Wang and Shu (2023), we sample 500 instances from the test set of each dataset using stratified sampling, ensuring a balanced label distribution. We use F1-score and accuracy as the primary evaluation metrics. To further assess generalizability, we also evaluate the accuracy on 500 positive [^2] CHECKWHY is a challenging benchmark. Its negative samples are created by modifying evidence to generate counterfactuals. Hence, we only use the positive samples. test samples from CHECKWHY (Si et al., 2024) in the main experiments. Details of the datasets can refer to Appendix A.1.

### v2

@@ Abstract
Multi-hop claim verification is inherently challenging, requiring multi-step reasoning to construct verification chains while iteratively searching for information to uncover hidden bridging facts. This process is fundamentally interleaved, as effective reasoning relies on dynamically retrieved evidence, while effective search demands reasoning to refine queries based on partial information. To achieve this, we propose Hierarchical Agent Reasoning and Information Search (HARIS), explicitly modeling the coordinated process of reasoning-driven searching and search-informed reasoning. HARIS consists of a high-level reasoning agent that focuses on constructing the main verification chain, generating factual questions when more information is needed, and a low-level search agent that iteratively retrieves more information, refining its search based on intermediate findings. This design allows each agent to specialize in its respective task, enhancing verification accuracy and interpretability. HARIS is trained using reinforcement learning with outcome-based rewards. Experimental results on the EX-FEVER and HOVER benchmarks demonstrate that HARIS achieves strong performance, greatly advancing multi-hop claim verification.  [^1] Code: https://github.com/qishenghu/HARIS
@@ 1 Introduction
Claim verification (Guo et al., 2022) has become a critical challenge as misinformation proliferates online. It requires systems to determine whether a given claim is supported or refuted based on retrieved evidence. While verifying simple claims involves shallow reasoning within a single document, the verification of complex, multi-hop claims presents a fundamentally different challenge. This difficulty stems from the fragmented nature of evidence (Pham et al., 2025; Atanasova et al., 2022). Effectively verifying such claims requires a joint process of multi-step reasoning and iterative information searching (Zheng et al., 2025).
Figure 1: Example of a challenging multi-hop verification. Verifying the claim demands coordinating reasoning-guided search to disambiguate entities and search-informed reasoning to adapt based on retrieved evidence. Prematurely concluding on any distracting branch (in red) leads to incorrect judgment. The correct path—from Baby (2007) to Overwatch via Feodor Chin—emerges only through this dynamic interplay, not static decomposition or single-pass retrieval.
A key challenge in multi-hop verification is identifying the correct bridging facts—implicit links that connect separate pieces of evidence but are not explicitly stated. As shown in Figure 1, verifying this claim requires identifying Feodor Chin as the critical bridging fact between the film Baby (2007) and the character Zenyatta from the game Overwatch, among other irrelevant paths. This process necessitates intensive interactions between reasoning, which proposes candidate hypotheses to identify potential bridging facts, and iterative search, which retrieves evidence to validate or eliminate certain hypotheses. In particular, reasoning is necessary to construct verification chains, but effective reasoning depends on relevant evidence, which often requires iterative search. Meanwhile, effective search relies on reasoning to formulate queries based on partially retrieved evidence. This reciprocal relationship, where reasoning shapes search and retrieved evidence refines ongoing reasoning, captures the recursive nature of multi-hop verification.
@@
To tackle these, we propose HARIS—Hierarchical Agent Reasoning and Information Search, explicitly modeling the coordinated process of reasoning-driven searching and search-informed reasoning. HARIS consists of two specialized large language model (LLM) agents: a high-level reasoning agent and a low-level search agent, both trained using reinforcement learning (RL) to optimize their respective tasks. The high-level agent forms the main verification chain, generating factual questions when more information is needed. The low-level agent handles these questions through dynamic search, iteratively refining its queries based on partial results to progressively build a comprehensive evidence base. This design allows each agent to specialize in its respective task, enhancing both verification accuracy and interpretability by clearly modeling the mutually reinforcing interaction between reasoning and information searching. To our knowledge, HARIS is among the first RL-based cooperative agent approach for claim verification.
@@
  • HARIS is trained end-to-end via Group Relative Policy Optimization with outcome-based rewards, directly optimizing task performance without intermediate supervision.
@@ 2.1 Claim Verification
Claim verification research has been increasingly focusing on enhancing transparency during decision making (Zeng and Gao, 2024; Chen et al., 2024; Hu et al., 2025a). QACheck (Pan et al., 2023a) reformulates verification as a progressive question-answering (QA) task, validating claims through step-wise questioning. Structural approaches (Jeon and Lee, 2025; Pham et al., 2025), such as ProgramFC (Pan et al., 2023b) and FOLK (Wang and Shu, 2023), which use symbolic reasoning or reasoning program to enforce systematic verification. The Decompose-Then-Verify (Wanner et al., 2024b, a; Jiang et al., 2024; Hu et al., 2025b; Lu et al., 2025) paradigm focus on breaking down complex claims into simpler sub-claims for independent validation. Agentic approaches have emerged as a promising direction (Zhao et al., 2024). LoCal (Ma et al., 2025) employs a prompt-driven multi-agent framework emphasizing causal consistency, and BiDeV (Liu et al., 2025) uses specialized agents to address vagueness and redundancy. In contrast, HARIS formulates claim verification as a cooperative process between reasoning and search agents, and trains both agents jointly via reinforcement learning.
@@ 2.2 Reasoning & Searching
Recent work has expanded the reasoning capabilities of LLMs by integrating search mechanisms (Xiong et al., 2025; Guan et al., 2025; Sun et al., 2025). Notable methods like Search-o1 (Li et al., 2025) incorporate dynamic search into reasoning frameworks, improving factual accuracy in open-domain and multi-hop reasoning tasks leveraging LLMs. Furthermore, RL methods like Group Relative Policy Optimization (GRPO) (Shao et al., 2024) have been used to incentivize search capabilities of LLMs (Gao et al., 2025), encouraging models to generate context-aware queries and effectively integrate retrieved information (Jin et al., 2025; Song et al., 2025; Chen et al., 2025; Qian et al., 2025; Wu et al., 2025). Collectively, these approaches demonstrate that RL-trained search integration can effectively improve performance on knowledge-intensive tasks.
@@ 3.1 Why Reasoning and Search Agents?
Meanwhile, the search and reasoning agents closely collaborate to enhance overall performance. Interactive generation allows the agents to iteratively guide each other’s outputs. Coordinated training alternately optimizes the agents to strengthen their collaboration, as shown in Figure 2.
@@ 3.2 High-level Reasoning Agent
The high-level reasoning agent constructs the main verification chain, coordinating multi-step reasoning and generating factual questions for the search agent when additional information is needed.
@@ 3.2.1 Reasoning Agent Rollout
Following prior work (Chen et al., 2025; Jin et al., 2025), the reasoning agent’s rollout process uses special tags to define question actions. Specifically, the tags <question> and </question> indicate that an action to call the search agent should be invoked. Upon detecting the </question> tag, the generation is paused and the enclosed content will be regarded as the factual question and sent to the search agent. The reported information from search agent is wrapped within <result> tags and appended to the sequence to enable continued rollout. The rollout ends when a final verification is derived, wrapped within <verification> tags. The prompt template is provided in Appendix C.1.
@@ 3.2.2 Reasoning Agent Reward
The reasoning agent is responsible for performing multi-step reasoning over the gathered evidence to verify the claim. It provides a binary verification decision, and the reward is the correctness of this decision. The overall reward $R_{\text{high}}$ is given by:
@@ 3.3.1 Training Data Synthesis
To train the search agent, it is crucial that the training data not only includes diverse question-answer pairs but also closely aligns with the type of the questions generated by the high-level agent. This alignment ensures that the search agent can effectively collaborate with the reasoning agent during multi-hop verification. To create such training set, we sample questions from the high-level agent rollouts, pair them with the corresponding ground-truth evidence, and use GPT-4o (OpenAI, 2024b) to generate pseudo ground-truth answers. To maintain data quality, we filter out pairs where GPT-4o outputs "none" as the answer, typically removing about 10% of the data. This process synthesize data that is both contextually relevant and closely matched to the reasoning patterns of the high-level agent. Synthesis details and examples can be found in Appendix A.4 and Table C.2.
@@ 3.3.2 Search Agent Rollout
The search agent rollout process is similar to the reasoning agent. When the generation process encounters a </search> tag, it pauses and extracts the content wrapped within the tags as the search query. This query is then used to perform top-k retrieval from the knowledge corpus. The retrieved text is wrapped in <result> and </result> tags and appended to the paused sequence, allowing the rollout to continue iteratively until the agent determines that sufficient information has been gathered. At this point, the agent reports the collected evidence within <report> and </report> tags, marking the completion of the rollout. The prompt template can be found in Appendix C.1.
@@ 3.3.3 Search Agent Reward
We use a combination of format and LLM-as-a-Judge approaches for search agent rewards. The LLM-as-a-Judge approach evaluates the quality the gathered information using an LLM, comparing the final output against the pseudo ground-truth answer [^2] We use GPT-4o-mini(OpenAI, 2024a) for this purpose, as it is cost-efficient.. The evaluation score is computed as:
@@
where ${a_{\text{pred}}}$ is the search agent’s final output and ${a_{\text{gt}}}$ is the pseudo ground-truth answer. This approach is less strict than exact match (EM) metrics, better aligning with the search agent’s goal of collecting comprehensive, contextually relevant information rather than just short, exact answers. The implementation can be found in Appendix A.5.1. Formally, the reward $R_{\text{low}}$ is given by:
@@
To assess the reliability of LLM-as-a-Judge, we recruited two annotators [^3] https://www.prolific.com/ for assessing held-out samples using the same criteria as LLM. The results showed strong consistency with human judgment (Cohen’s Kappa: 0.81; agreement: 93.3%). See Appendix D for complete human evaluation details.
@@ 3.5 Coordinated Agent Training
To enable effective collaboration between the high-level reasoning agent and the low-level search agent, we adopt a Coordinated Agent Training strategy. This process consists of two stages: an initial foundation stage, where both agents develop core reasoning, search, and formatting abilities; and a subsequent alternating training stage, where the agents iteratively refine their specialized skills through mutual interaction.
Given a training set $T$ of claim verification data, in the foundation stage, the untrained low-level search agent $L_{0}$ and high-level reasoning agent $H_{0}$ are trained sequentially to establish their foundational capabilities. First, we sample questions from $H_{0}$ to train $L_{0}$, producing the updated search agent $L_{1}$. This updated search agent then supports the training of $H_{0}$, producing the updated reasoning agent $H_{1}$. This stage establishes a foundation for both agents in reasoning, searching, and formatting.
In the alternating stage, we promote further coordination by repeatedly alternating training between the two agents. The dataset $T$ is divided into $N$ segments. For each segment, we sample questions from the current high-level agent $H_{i}$ to train the low-level agent, producing $L_{i+1}$. The updated low-level agent $L_{i+1}$ is then used to collaborate with the high-level agent, resulting in $H_{i+1}$. This alternating process continues for $N$ rounds, fostering mutual adaptation and ensuring the agents remain closely aligned throughout training.
@@
Input: Initial low-level agent $L_{0}$, high-level agent $H_{0}$, training set $T$
@@
Stage 1: Foundation Training
@@
Stage 2: Alternating Training
Divide $T$ into $N$ segments $\{T_{1},\ldots,T_{N}\}$
for $i=1$ to $N$ do
@@
RAG 58.9658.9658.96 59.2059.2059.20 56.5956.5956.59 56.6056.6056.60 55.0655.0655.06 55.2055.2055.20 68.8368.8368.83 69.0069.0069.00 64.4164.4164.41 64.8064.8064.80 56.4056.4056.40
Decomp-Verify 62.3962.3962.39 62.6062.6062.60 57.2757.2757.27 57.3157.3157.31 54.5554.5554.55 55.6055.6055.60 68.0368.0368.03 68.4068.4068.40 61.9561.9561.95 63.0063.0063.00 36.0036.0036.00
ProgramFC 66.8466.8466.84 66.8066.8066.80 55.3555.3555.35 56.8056.8056.80 48.3048.3048.30 52.6052.6052.60 71.2871.2871.28 71.6071.6071.60 60.3460.3460.34 62.4062.4062.40 21.6021.6021.60
BiDeV 64.5164.5164.51 65.0065.0065.00 57.5957.5957.59 58.6058.6058.60 54.9454.9454.94 57.0057.0057.00 67.7567.7567.75 67.8067.8067.80 61.5261.5261.52 62.0062.0062.00 24.6024.6024.60
QACheck 67.6067.6067.60 67.6067.6067.60 60.6060.6060.60 60.6060.6060.60 58.9158.9158.91 59.0059.0059.00 75.5275.5275.52 75.6075.6075.60 68.4268.4268.42 68.6068.6068.60 58.0058.0058.00
FOLK 67.2267.2267.22 67.6067.6067.60 59.8959.8959.89 61.2061.2061.20 50.9050.9050.90 55.2055.2055.20 75.5575.5575.55 75.8075.8075.80 67.2467.2467.24 68.4068.4068.40 50.7050.7050.70
Search-o1 68.7268.7268.72 69.0069.0069.00 59.3459.3459.34 59.8059.8059.80 54.9054.9054.90 56.6056.6056.60 77.4177.4177.41 77.8077.8077.80 72.0872.0872.08 72.8072.8072.80 52.4052.4052.40
@@ 4.1 Datasets
For training, we sample 7,200 examples from the combined EX-FEVER and HOVER training data. For evaluation, following Wang and Shu (2023), we sample 500 instances from the test set of each dataset using stratified sampling, ensuring a balanced label distribution. We use F1-score and accuracy as the primary evaluation metrics. To further assess generalizability, we also evaluate the accuracy on 500 positive [^4] CHECKWHY is a challenging benchmark. Its negative samples are created by modifying evidence to generate counterfactuals. Hence, we only use the positive samples. test samples from CHECKWHY (Si et al., 2024) in the main experiments. Details of the datasets can refer to Appendix A.1.